OpenStudy (anonymous):

State end-of-grade tests tend to follow a normal distribution. In 2008, the mean test score for 8th graders was 76.5 with a standard deviation of 5. For each part of the problem below, show your work including a sketch with corresponding area shown. A.If a student scored at the 97th percentile, what score would he or she have? B.What percent of students score higher than 92? C.What percent of students score between 80 and 92? plz help

OpenStudy (boldjon):

You need the normal distribution table to solve the problem 1. The Z for 97% is 1.88. Then you multiply 5*1.88=9.4. Add that to 76.5 to get 85.9 2. Z=(92-76.5)/5=3.1 P(Z>3.1)=0 3.P(80<X<92)=P(X<92)-P(X<80) Z=(80-76.5)/5=0.7 Area=0.2580 P(X<92)-P(X<80)=0.5-0.2580=0.242

OpenStudy (anonymous):

thanks man, I did the problem was making sure it was right!